OpenStudy (anonymous):

Triangle PQR is a right triangle. If PQ=8, what is PR? RQ?

OpenStudy (anonymous):

Depends on where the right angle is relative to the side PQ, and I think you need at least one more piece of information, a side length or another angle.

OpenStudy (anonymous):

The triangle is 30-60-90

OpenStudy (anonymous):

Yeah that was critical information lol :D

OpenStudy (anonymous):

Hahah, sorry, I will make sure to post everything I know from now on. Also, would it be easier to just post a picture of the triangle or image itself along with my question?

OpenStudy (anonymous):

That's a 1 : 2 : \(\sqrt{3}\) ratio triangle then

OpenStudy (anonymous):

And which side is PQ?

OpenStudy (anonymous):

Btw, welcome to Open Study :-D Da rulez: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

So PQ is your hypotenuse then...

OpenStudy (anonymous):

Yessir, thought it might be easier to see the picture.

OpenStudy (anonymous):

I would always recommend posting the pictures with geometry stuff :-D

OpenStudy (anonymous):

Haha yeah, seems much easier to understand the problem then, my problem is I always flip the picture around in my head and mess up the numbers

OpenStudy (anonymous):

Remember this? \[\huge a^2 + b^2 = c^2\] C = your PQ side, 8

OpenStudy (anonymous):